OpenStudy (frank0520):

Use the Divergence Theorem to calculate the flux of F(x,y,z)=(x^5 +y^5 +z^5 -2x-3y-4z)i+sin(2y)j+4z(sin(y))^2 k across the surface of the tetrahedron bounded by the coordinate planes and the plane x+y+z=1.

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

\[\iint\limits_{S}\vec{F}\bullet d\vec{S}~=~\iiint \limits_{E}\text{div}(\vec{F}) dV \]

ganeshie8 (ganeshie8):

work the divergence and plug it in

OpenStudy (anonymous):

we can write this as a volume integral over \(E=\{(x,y,z)\in\mathbb{R}^3:x+y+z\le 1, x\ge 0, y\ge0, z\ge0\}\) using the divergence theorem as ganeshie8 pointed out, which is something like: $$\iint\limits_S F\cdot dS=\iiint\limits_E\nabla\cdot F\ dV$$

OpenStudy (frank0520):

\[\int\limits_{0}^{1}\int\limits_{0}^{1-x}\int\limits_{1}^{1-x-y}dzdydx\]

OpenStudy (anonymous):

now, using Cartesian coordinates so \(dV=dz\ dy\ dx\) $$\int_0^1\int_0^{1-x}\int_0^{1-x-y}\nabla\cdot F\ dz\ dy\ dx$$ clearly \(\nabla\cdot F=(5x^4-2) +2\cos(2y)+4\sin^2(y)\) so: $$\int_0^1\int_0^{1-x}\int_0^{1-x-y}\left(5x^4-2+2\cos(2y)+4\sin^2(y)\right)\,dz\,dy\,dx$$

OpenStudy (frank0520):

can you tell me if I have the correct limits

OpenStudy (anonymous):

no, \(0\le z\le 1-x-y\)

OpenStudy (frank0520):

Thanks for the help, I had a typo with the dz limit